# Critical fragility in socio-technical systems

### v3

@@ Abstract
Optimizing for improving efficiencies can make socio-technical systems fragile. Many optimization problems in STSs belong to the class of constrained optimization: either through the efforts of system operators or by betting on the virtue of an invisible hand, such systems strive to achieve some kind of optimality, with a large amount of constraints. Indeed, optimizing cost and increasing service frequency constrained to reusing the crew from the incoming trains for outgoing trains, as well as input-output production networks — where each firm attempts to maximize profits by appropriately choosing its suppliers and production technologies to match demand from its customers while minimizing inventories — are real-world, and complex, examples of constrained optimization problems.
@@
As noted above, operators of STSs therefore have a natural tendency to perceive certain safeguards ensuring stability and resilience as unnecessary hindrances to achieving ever higher efficiency. Further, this tendency is particularly strong when the STS has displayed stability for a long time, making such safety measures or regulations seem unjustified. Driven by the desire to optimize, they come to dismantle these safeguards, inadvertently edging the system critically close to collapse. This scenario resembles that of Self-Organized Criticality (SOC) [33, 34, 35, 36] (for the case of STSs, it is organized by the operators themselves that are part of the systems), illustrated by a sand pile that keeps growing until its slope becomes so steep that an avalanche occurs, (partly) collapsing the sand pile. This is also the mechanism proposed by Minsky to explain the recurrence of financial crises [37]: regulations limiting the risk of financial institutions are put in place after crashes, thereby lowering the collective perception of risk. Once the memory of the last crisis fades, these safeguards are progressively removed, bringing about another crisis.

### v4

@@ Abstract
Optimizing for efficiencies can make socio-technical systems fragile. Many optimization problems in STSs belong to the class of constrained optimization: either through the efforts of system operators or by betting on the virtue of an invisible hand, such systems strive to achieve some kind of optimality, with a large amount of constraints. Indeed, optimizing cost and increasing service frequency constrained to reusing the crew from the incoming trains for outgoing trains, as well as input-output production networks — where each firm attempts to maximize profits by appropriately choosing its suppliers and production technologies to match demand from its customers while minimizing inventories — are real-world, and complex, examples of constrained optimization problems.
@@
In principle, faced with uncertainty about parameter values, system operators should adopt robust strategies. However, as noted above, operators of STSs have a natural tendency to perceive certain safeguards ensuring stability and resilience as unnecessary hindrances to achieving ever higher efficiency. This tendency is particularly strong when the STS has displayed stability for a long time, as overconfidence and/or complacency on the part of system operators make such safety measures or regulations seem unjustified. Driven by the desire to optimize, they come to dismantle these safeguards, inadvertently edging the system critically close to collapse. This is precisely the mechanism proposed by Minsky to explain the recurrence of financial crises [37]: regulations limiting the risk of financial institutions are put in place after crashes, thereby lowering the collective perception of risk. Once the memory of the last crisis fades, these safeguards are progressively removed, bringing about another crisis. Such a scenario resembles that of Self-Organized Criticality (SOC) [33, 34, 35, 36] (for the case of STSs, it is organized by the operators themselves that are part of the systems), illustrated by a sand pile that keeps growing until its slope becomes so steep that an avalanche occurs, (partly) collapsing the sand pile.